# NavMoE short=NavMoE, long=Navigation via Mixture of Experts

### v1

@@ III-B Proof of Convergence Bound
Figure 4: $\mathcal{T}$ is modeled as an undirected graph $G=(V,E)$. A path through the environment is modeled as a sequence of vertices $\tau=[v_{1},v_{2},\cdots,v_{t},\cdots,v_{T}]$ connected via $G$.
To support the proof, we model a traversability map $\mathcal{T}$ as an undirected graph $G=(V,E)$, where each vertex $v\in V$ corresponds to a pixel in $\mathcal{T}$ that is assigned a scalar traversability value (as shown in Fig. 4). An edge $(v_{i},v_{j})\in E$ exists if the pixels are adjacent, capturing local connectivity. A path through the environment is defined as a sequence of vertices $\tau=[v_{1},v_{2},\cdots,v_{t},\cdots,v_{T}]$ connected via $G$. The path planner seeks an optimal path $\tau^{*}$ by minimizing an objective function subject to feasibility constraints. We formulate the path planning problem within graph $G$ as a constrained optimization problem $\mathcal{P}$:
@@
From the definitions of $\underline{\mathcal{T}_{k}}$ and $\overline{\mathcal{T}_{k}}$ in Eqs. 9a and 9b, it follows that for any vertex $v_{t}$ along a path $\tau$, we have:
@@ III-B2 Monotonicity of $\delta_{k}$
Similarly, for the pessimistic estimate:
$\small\underline{\mathcal{T}_{k}}-\underline{\mathcal{T}_{k+1}}=-\hat{\mathcal{T}}_{k+1}\cdot\frac{\mathbf{W}_{k+1}}{\sum_{j=1}^{K}\mathbf{W}_{j}}\leq 0\Rightarrow\underline{\mathcal{T}_{k+1}}\geq\underline{\mathcal{T}_{k}}$ (21)
@@ IV-A Pre-Training on Large-Scale Partially-Aligned Data
TABLE I: Semantics to assigned expert. M: model-based method; NN: Neural Networks.
@@ V Experiment Results
To evaluate the cross-domain traversability estimation ability and computational efficiency of NavMoE, we compare our approach with 4 state-of-the-art baselines that can run in real-time: Falco [6], Mask2Former [25] for indoor, MaskFormer [26] for structured outdoor, and GANav [12] for unstructured outdoor environment. Additionally, we include four ablation studies, NavMoE without fine-tuning and lazy gating (Ours w/o FT/LG), without lazy gating (Ours w/o LG), without lower-level lazy gating (Ours w/o LLG) and without fine-tuning (Ours w/o FT), to evaluate the individual contributions of these components to overall performance. Scene overview for all three domains are shown in Fig. 5.A. The indoor experiment is conducted in real world, outdoor experiments use test sequences from [24, 27].
We evaluate performance using four metrics: Motion Primitives Quality: We introduce aligned-path traversability estimation mean square error ($e_{P}$) to measure the mean squared error between estimated and ground-truth traversability along candidate motion primitives, which directly reflects planning quality with arbitrary goals. It can be formulated as: $e_{P}=\left\lVert S(\mathcal{T})-S(\mathcal{T}_{gt})\right\rVert_{2}^{2}\big{/}\left|S(\mathcal{T}_{gt})\right|$. Traversability Map Quality: We introduce traversability map estimation mean square error ($e_{M}$) to assess the accuracy of the predicted traversability map; Computational Efficiency: is quantified in GFLOPS; and Normalized path quality ($Q_{p}$): Given any path $\tau$, goal $p_{goal}$, and $\mathcal{T}$, we define the path quality as $\eta(\tau)=1-f(\tau;\mathcal{T}_{gt},p_{goal})$, where $f(\tau;\mathcal{T},p_{goal})\in[0,1]$ denotes the path cost (as defined in Eq. 4). Path quality ratio $Q_{p}$ is then formulated as: $Q_{p}=\eta(\tau_{est})/\eta(\tau_{gt})$. This metric ranges between 0 and 1, with higher values indicating better path quality. Since our approach involves the lazy gating mechanism that doesn’t need to produce a full traversability map for path planning, we don’t record $e_{P}$ and $e_{M}$ for our approach.
We report GFLOPs instead of runtime because some experts in our MoE framework are implemented in C++ while others in Python. Thus, we believe GFLOPs provide a fairer metric for comparison. We here report the average runtimes estimated in outdoor scenario to demonstrate our time efficiency: Gating network 6.0ms; Model-based expert 6.2ms; Learning-based expert 64.3ms; and path calculation 119.8ms, and the average processing time is 163.8ms. We believe this is good for common Navigation, although the code is not optimized yet.
@@
Quantitative Results As illustrated in Fig. 5.B, our method navigates complex environments, such as Fig. 5.B 1, safely by primarily leveraging model-based experts while querying learning-based experts only when semantic cues are critical, achieving robust performance across all domains and terrain types. In Fig. 5.B 3, learning-based expert [26] fail to strictly adhere to safety constraints, potentially causing the robot falling off a curb into traffic, while NavMoE utilizes geometric perception from model-based methods to maintain safety. Conversely, in Fig. 5.B 2 and 4, the learning-based experts’ semantic perception correctly identifies non- or low-traversable areas (e.g., stop sign, grass), whereas the model-based method [6] mistakenly classify grass as traversable due to its geometric similarity to planar surfaces.
@@
Cross-Domain Traversability Estimation Accuracy. As shown in Table II, NavMoE is the only method that can generalize well across all three domains, achieving best or second best performance among $e_{P}$, $e_{M}$ and $Q_{p}$. In structured outdoor environment where learning-based expert [26] and Model-based one [6] have complementary expertise, the advantage of NavMoE becomes significant and we achieve the best performance among $e_{P}$, $e_{M}$ and $Q_{p}$. In indoor environment, our approach achieves comparable accuracy compared to domain-specific baseline [25] with only a $0.03$ drop in $Q_{p}$. The same phenomenon extends to unstructured outdoor environment with a $0.01$ drop in $Q_{p}$. This trend aligns with findings reported in other MoE works [16, 28], which indicates that MoE methods tend to deliver similar or occasionally slightly lower performance in domains where a single expert already excels. Additionally, our results are impacted by hardware misalignment and synchronization issues between the RGB camera and lidar or depth camera, which can be ovserved in Model-based method part in Fig. 5. Effectiveness of FT: Comparing NavMoE w/o FT/LG to NavMoE w/o LG, fine-tuning improves estimation accuracy up to $41\%$ in $e_{P}$ and $54\%$ in $e_{M}$. When lazy gating is applied (Ours w/o FT v.s. Ours in Table II), fine-tuning balances $Q_{p}$ and GFLOPS. Specifically, in the indoor scenario, fine-tuning shifts weights toward model-based methods, improving efficiency with small $Q_{p}$ loss. Conversely, in outdoor environments, fine-tuning allocates more weights to learning-based methods, enhancing $Q_{p}$.
@@
Efficiency and Effectiveness for Path Planning. Across all domains, NavMoE has substantially higher efficiency than the best performing expert, marked in blue in Table II, by up to $76\%$ With the guarantee solution quality introduced by the lazy gating mechanism, The result $Q_{p}$ are within $0.03$ compared to NavMoE w/o LG, which is within the bound set as $0.05$. Effectiveness of LG: Comparing NavMoE to our method w/o LG, lazy gating can decrease the computational overhead up to $96.3\%$, with only up to $0.02$ loss in $Q_{p}$. Interestingly, we found the upper level lazy router achieves over $95\%$ confidence and over $98\%$ in accuracy across all domains, meaning that keeping the domain lazy router will not affect the estimation accuracy but significantly improve the computational efficiency, shown in Table II: Ours w/o LG and Ours w/o LLG. In this case, our method can still improve the computational efficiency by up to $82.6\%$, thereby reducing the average inference time to 164 milliseconds.

### v2

@@ III-B Proof of Convergence Bound
To support the proof, we model a traversability map $\mathcal{T}$ as an undirected graph $G=(V,E)$, where each vertex $v\in V$ corresponds to a pixel in $\mathcal{T}$ that is assigned a scalar traversability value. An edge $(v_{i},v_{j})\in E$ exists if the pixels are adjacent, capturing local connectivity. A path through the environment is defined as a sequence of vertices $\tau=[v_{1},v_{2},\cdots,v_{t},\cdots,v_{T}]$ connected via $G$. The path planner seeks an optimal path $\tau^{*}$ by minimizing an objective function subject to feasibility constraints. We formulate the path planning problem within graph $G$ as a constrained optimization problem $\mathcal{P}$:
@@
From the definitions of $\underline{\mathcal{T}_{k}}$ and $\overline{\mathcal{T}_{k}}$ in Eqs. 7a and 7b, it follows that for any vertex $v_{t}$ along a path $\tau$, we have:
@@ III-B2 Monotonicity of $\delta_{k}$
Similarly, for the pessimistic estimate we have $\underline{\mathcal{T}_{k+1}}\geq\underline{\mathcal{T}_{k}}$.
@@ IV-A Pre-Training on Large-Scale Partially-Aligned Data
TABLE I: Semantics to assigned expert. M: model-based method; NN: Neural Networks. The definition of traversability cost $c$ follows the prior work [21].
@@ V Experiment Results
To evaluate the cross-domain traversability estimation ability and computational efficiency of NavMoE, we compare our approach with 4 state-of-the-art baselines that can run in real-time: Falco [6], Mask2Former [26] for indoor, MaskFormer [27] for structured outdoor, and GANav [12] for unstructured outdoor environment. Additionally, we include four ablation studies, NavMoE without fine-tuning and lazy gating (Ours w/o FT/LG), without lazy gating (Ours w/o LG), without lower-level lazy gating (Ours w/o LLG) and without fine-tuning (Ours w/o FT), to evaluate the individual contributions of these components to overall performance. Scene overviews for all three domains are shown in Fig. 5.A. The indoor experiment is conducted in the real world, while outdoor experiments use test sequences from [25, 28].
We evaluate performance using four metrics: Motion Primitives Quality: We introduce aligned-path traversability estimation mean square error ($e_{P}$) to measure the mean squared error between estimated $\mathcal{T}$ and ground-truth traversability $\mathcal{T}_{\textrm{gt}}$ along candidate motion primitives, which directly reflects planning quality with arbitrary goals. It can be formulated as: $e_{P}=\left\lVert S(\mathcal{T})-S(\mathcal{T}_{\textrm{gt}})\right\rVert_{2}^{2}\big{/}\left|S(\mathcal{T}_{\textrm{gt}})\right|$. Traversability Map Quality: We introduce traversability map estimation mean square error ($e_{M}$) to assess the accuracy of the predicted traversability map; Computational Efficiency: is quantified in GFLOPS; and Normalized path quality ($Q_{p}$): Given any path $\tau$, goal $p_{\textrm{goal}}$, and $\mathcal{T}$, we define the path quality as $\eta(\tau)=1-f(\mathcal{T},p_{\mathrm{goal}})$, where $f(\mathcal{T},p_{\mathrm{goal}})\in[0,1]$ is defined in Eq. 4). The path quality ratio $Q_{p}$ is then formulated as: $Q_{p}=\eta(\hat{\tau})/\eta(\tau_{\textrm{gt}})$, where $\tau_{\textrm{gt}}$ represents the optimal ground truth path and $\hat{\tau}$ means the estimated path. This metric ranges between 0 and 1, with higher values indicating better path quality. Since our approach involves the lazy gating mechanism that doesn’t need to produce a full $\mathcal{T}$ for path planning, we do not record $e_{P}$ and $e_{M}$ for our approach.
We report GFLOPs instead of runtime because some experts in our MoE framework are implemented in C++ while others in Python. Thus, we believe GFLOPs provide a fairer metric for comparison. We here report the average runtimes estimated in outdoor scenario to demonstrate our time efficiency: Gating network 6.0ms; Model-based expert 6.2ms; Learning-based expert 64.3ms; and path calculation 119.8ms, and the average processing time is 163.8ms. We believe this is sufficient for common navigation tasks, although the code implementation is not optimized yet.
@@
Quantitative Results.  As illustrated in Fig. 5.B, our method navigates complex environments, such as Fig. 5.B 1, safely by primarily leveraging model-based experts while querying learning-based experts only when semantic cues are critical, achieving robust performance across all domains and terrain types. In Fig. 5.B 3, learning-based expert [27] fail to strictly adhere to safety constraints, potentially causing the robot falling off a curb into traffic, while NavMoE utilizes geometric perception from model-based methods to maintain safety. Conversely, in Fig. 5.B 2 and 4, the learning-based experts’ semantic perception correctly identifies non- or low-traversable areas (e.g., stop sign, grass), whereas the model-based method [6] mistakenly classify grass as traversable due to its geometric similarity to planar surfaces.
@@
Cross-Domain Traversability Estimation Accuracy.  As shown in Table II, NavMoE is the only method that can generalize well across all three domains, achieving best or second best performance among $e_{P}$, $e_{M}$ and $Q_{p}$. In structured outdoor environment where learning-based expert [27] and Model-based one [6] have complementary expertise, the advantage of NavMoE becomes significant and we achieve the best performance among $e_{P}$, $e_{M}$ and $Q_{p}$. In indoor environment, our approach achieves comparable accuracy compared to domain-specific baseline [26] with only a $0.03$ drop in $Q_{p}$. The same phenomenon extends to unstructured outdoor environment with a $0.01$ drop in $Q_{p}$. This trend aligns with findings reported in other MoE works [16, 29], which indicates that MoE methods tend to deliver similar or occasionally slightly lower performance in domains where a single expert already excels. Additionally, our results are impacted by hardware misalignment and synchronization issues between the RGB camera and lidar or depth camera, which can be observed in the model-based method part in Fig. 5. Effectiveness of FT.  Comparing NavMoE w/o FT/LG to NavMoE w/o LG, fine-tuning improves estimation accuracy up to $41\%$ in $e_{P}$ and $54\%$ in $e_{M}$. When lazy gating is applied (Ours w/o FT v.s. Ours in Table II), fine-tuning balances $Q_{p}$ and GFLOPS. Specifically, in the indoor scenario, fine-tuning shifts weights toward model-based methods, improving efficiency with small $Q_{p}$ loss. Conversely, in outdoor environments, fine-tuning allocates more weights to learning-based methods, enhancing $Q_{p}$.
Efficiency and Effectiveness for Path Planning.  Across all domains, NavMoE has much higher efficiency than the best performing expert, marked in blue in Table II, by up to $76\%$. With the guaranteed solution quality introduced by the lazy gating mechanism, the resulting $Q_{p}$ are within $0.03$ compared to NavMoE without LG, which is within the bound set as $0.05$.
@@
Effectiveness of LG.  Comparing NavMoE to our method without LG, lazy gating can decrease the computational overhead up to $96.3\%$, with only up to $0.02$ loss in $Q_{p}$. Interestingly, we found the upper level lazy router achieves over $95\%$ confidence and over $98\%$ in accuracy across all domains, meaning that keeping the domain lazy router will not affect the estimation accuracy but significantly improve the computational efficiency, shown in Table II: Ours w/o LG and Ours w/o LLG. In this case, our method can still improve the computational efficiency by up to $82.6\%$, thereby reducing the average inference time to 164 milliseconds.